OpenStudy (anonymous):

Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. Help Please!

OpenStudy (anonymous):

9-50-81+81 = -59

OpenStudy (anonymous):

okay Thank You!

OpenStudy (anonymous):

oh no , is wrong :) \[3^{3} - 2X 5^{2} - 3 X 3^{3} + (-3)^{4} = 27 - 50 - 81 + 81 = -23 :D\]

OpenStudy (anonymous):

Okay thank you. I get it now!